To calculate the annual loss in the mass of the sun if approximately 8.4J of radiated energy is received by each square cm of the earth's surface per minute. The distance of the sun from the earth's surface is 1.15×10^11m.
1
Expert's answer
2019-12-27T08:57:19-0500

The energy radiated by the sun is"W = IT4\pi R^2 = 8.4\cdot60\cdot24\cdot30\cdot12\cdot10^4\cdot4\cdot3.14\cdot(1.15\cdot10^{11})^2 = 7.2\cdot10^{33} J."

This is equivalent to annual weight loss "m = \frac{W}{c^2} = \cfrac{7.2\cdot10^{33}}{9\cdot10^{16}} = 0.8\cdot10^{17} kg."